# Online Process Reward Leanring for Agentic Reinforcement Learning

### v2

@@ 3.2 Theoretical Analysis
We now justify that the learned step reward $r_{\phi}$ defined in Eq. 1 is (1) an unbiased estimator of an action’s contribution to the latent task utility, and (2) a safe, well-conditioned learning signal for the policy $\pi_{\theta}$. We start from a Bradley–Terry (BT) preference (Bradley & Terry, 1952) assumption on pairwise trajectories.
Assumption 3.2 (Trajectory-level preference model).
There exists a latent utility function $R^{\star}:\tau\!\mapsto\!\mathbb{R}$ such that the outcome reward verifier or model chooses $\tau^{+}$ over $\tau^{-}$ with probability
$\Pr(\tau^{+}\succ\tau^{-})\;=\;\sigma\!\Bigl(\tfrac{1}{\beta}\bigl[R^{\star}(\tau^{+})-R^{\star}(\tau^{-})\bigr]\Bigr).$
This is the classical BT likelihood for pairwise comparisons. Minimizing the PRM loss in Eq. 2 is therefore exact maximum-likelihood estimation of a score $\log\pi_{\phi}(\tau)-\log\pi_{\theta_{old}}(\tau)$ that matches $\frac{R^{\star}(\tau)}{\beta}$ up to a constant.
Theorem 3.3 (Potential-based reward shaping).
Let the PRM $\pi_{\phi}$ be a global optimizer of the loss in Eq. 2, and let $\pi_{ref}$ be the static reference model used in a DPO loss. Then $r_{\phi}(\tau)\;=\;\log\pi_{\phi}(\tau)\;-\;\log\pi_{\text{ref}}(\tau)=\tfrac{1}{\beta}\,R^{\star}(\tau)+c(x)$ for some prompt-dependent constant $c(x)$. With the step-reward definition (Eq. 1) we obtain the telescoping identity:
$\sum_{t=1}^{T}r_{\phi}(o_{1:t},a_{t})=\;R^{\star}(\tau)+\beta c(x)+\beta\!\Bigl[\log\pi_{\text{ref}}(\tau)-\log\pi_{\theta_{\text{old}}}(\tau)\Bigr].$
Choosing $\theta_{\text{old}}=\theta$ deletes the last term. Therefore, $r_{\phi}$ is a potential-based shaping of $R^{\star}$ that preserves the optimal policy set (Ng, 2003).
Proposition 3.4 (Gradient alignment and boundedness).
Define the shaped objective
$J(\theta)\;=\;\mathbb{E}_{\tau\sim\pi_{\theta}}\!\Bigl[\sum_{t=1}^{T}r_{\phi}(o_{1:t},a_{t})\Bigr].$
Then the policy gradient is:
$\nabla_{\theta}J(\theta)=\beta\,\mathbb{E}_{a_{t}\sim\pi_{\theta}}\!\bigl[\nabla_{\theta}\log\pi_{\theta}(a_{t}\!\mid\!o_{1:t})\,\bigl(\log\pi_{\phi}-\log\pi_{\theta_{\text{old}}}\bigr)\bigr].$
If $\theta_{\text{old}}=\theta$, this reduces to $\nabla_{\theta}J(\theta)=-\beta\,\nabla_{\theta}\,\mathrm{KL}\!\bigl(\pi_{\theta}\,\|\,\pi_{\phi}\bigr)$, so maximizing $J$ performs a KL-descent step that moves $\pi_{\theta}$ towards the preference-aligned PRM $\pi_{\phi}$ inside a trust region. Let $\varepsilon=\min_{a_{t},o_{1:t}}\pi_{\theta_{\text{old}}}(a_{t}\!\mid\!o_{1:t},x)>0$. Then
$|r_{\phi}(o_{1:t},a_{t})|\leq\beta\log\bigl(1/\varepsilon\bigr).$
Hence policy-gradient samples are uniformly bounded, guaranteeing stable stochastic optimization of the alternating $\{\pi_{\phi},\pi_{\theta}\}$ updates.
In summary, since $\pi_{\phi}$ concentrates probability on behaviors that increase $R^{\star}$, each $r_{\phi}$ offers an approximately unbiased estimate of an action’s contribution to the latent task utility. The dense, bounded step rewards also reduces variance relative to a single outcome reward, thus yielding an effective credit assignment strategy to accelerate policy learning.
@@ 4 Experiments
We evaluate OPRL across a variety of agentic tasks to demonstrate: (1) its effectiveness in training LLM agents for long-horizon reasoning and acting; (2) high sample efficiency and training stability given step rewards learned online; (3) improved exploration justified by consistent reward increases; and (4) the core components of OPRL for credit assignment in agentic RL. Training details can be found in Appendix C.
@@
We evaluate LLM agents in three challenging environments: (1) WebShop (Yao et al., 2022), a text-based web environment where the agent interacts with a HTML-based website to search, nevigate, and purchase an item given an user instruction, requiring multi-step decision making; (2) VisualSokoban (Schrader, 2018) with $6\times 6$ size, a transportation puzzle game where the agent has to push all boxes on targets, requiring spatial reasoning and long-term planning over both visual and textual inputs; (3) SOTOPIA (Zhou et al., 2024), an open-ended role-playing environment where the agent interacts with another player given a social scenario, role profiles and private goals, requiring strategic reasoning over opponent’s real-time strategies. For training, we use scenarios from SOTOPIA-$\pi$ (Wang et al., 2024b). See Appendix B for more details of the environments.
@@
We compare our method against a range of competitive baselines: (1) Prompting LLMs specialized in general-purpose reasoning: GPT-5 (OpenAI, 2025b), Gemini-2.5-Pro (Google, 2024), DeepSeek-R1 (Guo et al., 2025), and Claude-Sonnet-4-Thinking (Anthropic, 2025); (2) RL post-training methods: PPO (Schulman et al., 2017), an actor-critic algorithm that requires an additional value model, and critic-free algorithms including GRPO (Shao et al., 2024), RLOO (Ahmadian et al., 2024) and REINFORCE++ (Hu et al., 2025). These RL approaches are used for trajectory-level optimization following Wang et al. (2025); (3) A recent single-turn RL method: PRIME (Cui et al., 2025) that introduces an online PRM to provide token-level process rewards for policy optimization; and (4) A recent agentic RL algorithm: GiGPO (Feng et al., 2025) that calculates step-level advantages via same-state grouping.
@@
All methods are evaluated using ReAct (Yao et al., 2023) prompting strategies, with CoT generated before each action. For WebShop and VisualSokoban, we follow Feng et al. (2025) to set the temparature of LLM agents to 0.4, and adopt the Success Rate and Score (only for WebShop) as the evaluation metrics, which are computed over validation instances and select the best score for comparison. For SOTOPIA, we report the goal completion score ranging from 0 to 10, which is evaluated by GPT-4o as a proxy for human judgement following Zhou et al. (2024). We set the tempature to 0 for the LLM judge and 0.7 for LLM agents.
Implementation details.
We use Qwen2.5-(VL)-7B-Instruct (Yang et al., 2024) and Llama3.1-8B-Instruct (Meta, 2024) as the base models for policy optimization. Our online PRM is initialized from the base policy model, except for VisualSokoban, where the policy model uses Qwen2.5-VL-7B-Instruct and the PRM uses Qwen2.5-7B-Instruct. We set the advantage coefficient $\alpha$ to 1 and $\beta=0.05$ for PRM training. The rollout size is set to 8 per prompt except PPO. We use RLOO as the default RL algorithm for policy update in WebShop and VisualSokoban, and use GRPO in SOTOPIA unless otherwise specified. See Appendix C for full training details.
4.1 Main Results
Performance on benchmarks.
Table 1 showcases OPRL’ superior performance over baselines in WebShop and VisualSokoban, with notable gains on the latter, where RL algorithms struggle with irreversible mistakes and limited foresight. Similar trends occur with much smaller base models as shown in Table 4. Specifically, OPRL surpasses recent multi-turn RL method GiGPO by enabling finer-grained credit assignment, distinguishing good and bad actions at each step with process rewards than relying solely on same-state grouping. Our method also outperforms PRIME by aligning turn-level optimization with step rewards instead of token-level rewards, simplifying policy learning and enhancing training stability. In SOTOPIA, where GiGPO and PRIME are inapplicable due to open-ended state space and unverifiable rewards, Table 2 shows that OPRL still achieves superior performance. Particularly, compared to RL post-training baselines, our method improves goal completion rate in hard scenarios by 14% under the self-chat setting, and up to 48% ($6.68\rightarrow 7.16$) increase when chatting with GPT-4o. This demonstrates the robustness and generalizability of our method to a wide variety of interactive environments.
Table 1: Performance on WebShop and VisualSokoban. Results are averaged over three random seeds. Qwen2.5-7B-Instruct and Qwen2.5-VL-7B-Instruct serve the base models for WebShop and VisualSokoban, respectively. Note that Deepseek-R1 and PPO training do not currently support multi-modal scenarios, and PRIME is only applicable to tasks with binary outcome rewards.
@@ 4.1 Main Results
+ OPRL $\textbf{86.5}\pm 2.8$ $\textbf{93.6}\pm 1.0$ $\textbf{91.7}\pm 1.2$
@@
+ OPRL $\textbf{7.11}\pm 0.19$ $\textbf{8.42}\pm 0.03$ $\textbf{6.76}\pm 0.18$ $\textbf{8.36}\pm 0.03$
@@
+ OPRL $\textbf{8.06}\pm 0.11$ $\textbf{9.20}\pm 0.03$ $\textbf{7.16}\pm 0.14$ $\textbf{8.45}\pm 0.03$
OPRL with different vanilla RL algorithms.
Since OPRL is compatible with any policy update method, we also evaluate it on different RL algorithms for LLM post-training, including RLOO, REINFORCE++ and GRPO, and compare each to its vanilla version that uses only episode rewards. As shown in Figure 3, OPRL consistently improves vanilla RL algorithms by integrating online step rewards for long-horizon credit assignment. For example, OPRL with RLOO obtains substantial gains of 6.3% in the success rate on both WebShop and VisualSokoban. Similar trends occur with REINFORCE++ and GRPO, demonstrating the robustness and adaptability to diverse RL algorithms and environments.
Figure 3: Performance comparison of OPRL with different vanilla RL algorithms. Results are reported using one seed. Qwen2.5-7B-Instruct and Qwen2.5-VL-7B-Instruct serve the base models for WebShop and VisualSokoban, respectively.
@@ 4.2 In-depth Analysis
Figure 4 illustrates that compared to PRIME, GiGPO, and vanilla RLOO, OPRL achieves faster improvement and higher final performance in validation metrics during RL training, demonstrating superior sample efficiency and training stability for interactive tasks with long-term dependencies. As shown in Figure 4(a), OPRL achieves the score of vanilla RLOO in WebShop in just 105 steps, around $2\times$ improvement in training efficiency. By 165 steps, our method reaches the highest score of 94.7%, demonstrating significant gains in both efficiency and performance. OPRL also achieves higher success rates earlier during RL training across environments. Notably, while PRIME exhibits comparable early-stage performance in WebShop, its growth stagnates and experiences sharp fluctuations. This is because overly fine-grained process rewards (e.g., token-level in PRIME) will complicate policy learning in multi-turn interactions, which usually involve much longer sequences than single-turn QA tasks. In contrast, OPRL continues to improve steadily, validating our algorithm design that a well-shaped step reward lowers gradient variance, producing smoother policy updates.
Figure 4: Validation performance during RL training in WebShop and VisualSokoban.
Figure 5: Training dynamics of OPRL in WebShop and VisualSokoban. Left: Dynamics of the episode and implicit step rewards during RL training by our method. Right: The episode length versus training step compared to baselines.
Reward signals and exploration efficiency.
To demonstrate the implicit step rewards provides useful feedback for agents, we visualize the dynamics of both episode- and step-reward in Figure 5(a) and (b). It can be seen that the step rewards improve very early (especially in VisualSokoban) and then episode reward follows, indicating that OPRL first captures good local action heuristics and then composes them into higher-return trajectories. As a result, it enables OPRL to reduce unnecessary actions during interactions, leading to shorter episode lengths. As shown in Figure 5(c) and (d), episode lengths decrease without compromising task success, as evidenced by the consistent increase in episode-rewards.
Table 3: Ablation studies on core components of OPRL. “RLOO”: only outcome rewards are used to compute advantages for policy update. “w/ ground-truth PR”: use raw step rewards provided by VisualSokoban to calculate step-level advantages. “w/ merged rewards”: implicit step rewards are added directly to outcome rewards before advantage computation. “w/ token-level PR”: implicit step rewards are computed at token level rather than step level. Results are reported using one seed.
@@
w/ ground-truth PR - - 87.5
@@
w/ token-level PR 82.0 90.0 89.1
w/ OPRL 89.1 94.7 93.0
@@ 4.3 Ablation Studies
Table 3 presents results of ablation experiments to validate the necessity of OPRL’ key components for effective credit assignment. First, raw step penalties provided in VisualSokoban show limited improvement over vanilla RL, suggesting that the step rewards learned by OPRL are superior credit signals for LLM agents. Second, merging step rewards into episode rewards obtains gains over vanilla RL but the improvement are modest compared to OPRL. This indicates that we should not only reward intermediate actions but also gate credit by final task success to prevent speculative reward exploitation. Therefore, combining signals at the advantage level is crucial to credit assignment. Third, learning token-level process rewards is sub-optimal to multi-turn RL, suggesting that overly fine-grained rewards may introduce noise and thus increase difficulty of policy learning.
@@ 5 Related Work
PRMs has been widely explored in single-turn problems, such as mathematical reasoning and single-shot code generation. In these settings, PRMs are used to score intermediate CoT steps for test-time search or reranking (Lightman et al., 2024; Wang et al., 2024a; Mahan et al., 2024), or for online-RL training (Dou et al., 2024; Setlur et al., 2025; Zha et al., 2025). Moving to interactive tasks, process rewards are usually constructed in three ways: (1) rule-based verifiable signals that assign turn-level rewards to tool execution (Zeng et al., 2025) or meta-reasoning behavior tags (Zhang et al., 2025b); (2) GRMs (Liu et al., 2025; Zou et al., 2025) that label step quality; (3) and implicit PRMs (Yuan et al., 2025) trained from outcome labels and converted into token-level rewards.
@@
However, manually-designed or judge-based step labels are costly and biased, suffering from reward hacking. Learning step Q-values (Choudhury, 2025) reduces the bias but a fixed PRM may poorly estimate Q-values for unseen actions during inference. PRIME(Cui et al., 2025) partially addresses these issues by jointly training an implicit PRM with the generator. Yet with overly fine-grained rewards in token-level, it introduces noises and distabilizes multi-turn RL training. Instead of learning a PRM, Feng et al. (2025) addresses credit assignment by computing step-level advantages via same-state grouping. While effective in tasks with finite state-action space, such a method hinges on exact state overlaps with limited generalization in open-ended language environments. In contrast to these methods, OPRL maintains an online PRM trained from trajectory-level preferences and converts it into dense step-level guidance, offering scalable, label-efficient step rewards with low-variance and showing robustness and generalization to both verifiable and unverifiable rewards in open-ended environments.
6 Conclusion and Limitations
We propose OPRL, a general-purpose credit–assignment strategy for long-horizon LLM agents. OPRL trains an online PRM, reusing the same on-policy rollouts that are generated to update the agent’s policy. A DPO-style objective converts trajectory-level preferences into dense, step-level rewards, which are then fused with episode-level advantages for policy update. Theoretically, the resulting step rewards are preference-consistent, potential-based shaping and bounded, ensuring low-variance gradient updates to stablize multi-turn training. In practice, OPRL can be plugged into standard RL algorithms and generalized to both verifiable and unverifiable rewards in open-ended settings. Empirical results show that OPRL achieves averaged 86.5% success and 93.6% score on WebShop, and reaches 91.7% success on VisualSokoban. In SOTOPIA, OPRL improves goal completion in hard scenarios by up to 14% in self-chat and 48% against GPT-4o.
While OPRL significantly advances credit assignment for agentic RL, several limitations remain. First, despite the use of our PRM for online RL, it can also be applied to test-time scaling for search guidance. Second, in SOTOPIA, the PRM is trained solely to predict goal-completion preferences, and future work cluld extend OPRL to multi-objective PRMs. Finally, the PRM and the policy can share one backbone with two heads—or a unified model trained with different objectives, thus reducing computation cost and potentially improving representation sharing.

### v3

@@ 3.2 Theoretical analysis
We now justify that the learning objective of our implicit PRM in Eq. 2 equivalents to a BT model with a step-wise reward function. Formally, for any trajectory pair $\{\tau_{i}=\{(o_{t}^{i},a_{t}^{i})\}_{t=1}^{T}\}_{i=1}^{2}$ satisfying $o_{1}^{1}=o_{1}^{2}$ (the same task prompt $x$), we have
$\begin{split}&\mathbb{P}(\tau_{1}\succ\tau_{2})=\sigma\!\left(\sum_{t=1}^{T}\beta\log\frac{\pi^{*}_{\phi}(a_{t}^{1}\mid o_{1:t}^{1},x)}{\pi_{\theta_{\mathrm{old}}}(a_{t}^{1}\mid o_{1:t}^{1},x)}-\sum_{t=1}^{T}\beta\log\frac{\pi^{*}_{\phi}(a_{t}^{2}\mid o_{1:t}^{2},x)}{\pi_{\theta_{\mathrm{old}}}(a_{t}^{2}\mid o_{1:t}^{2},x)}\right)\\ &=\sigma\!\left(\sum_{t=1}^{T}r_{\phi}^{*}(o_{1:t}^{1},a_{t}^{1})-\sum_{t=1}^{T}r_{\phi}^{*}(o_{1:t}^{2},a_{t}^{2})\right)\end{split}$ (7)
where $\pi_{\phi}^{*}$ denotes the optimal implicit PRM. Based on the autoregressive nature of policies, Eq. 7 aligns with the learning objective of DPO (Rafailov et al., 2024b), but with $\pi_{\theta_{\text{old}}}$ as the reference model. Similar to the learning objective in a token-wise setup in Zhong et al. (2025) that is equivalent to a BT model with a token-wise reward function $r^{*}(x,y)=\beta\log\frac{\pi_{\phi}^{*}(y_{t}|x,y_{1:t-1})}{\pi_{\text{ref}}(y_{t}|x,y_{1:t-1})}$, Eq. 7 shows that the learning objective in Eq. 2 equivalents to a BT model with a step-wise reward function:
$r_{\phi}^{*}(o_{1:t},a_{t})=\beta\log\frac{\pi^{*}_{\phi}(a_{t}\mid o_{1:t},x)}{\pi_{\theta_{\mathrm{old}}}(a_{t}\mid o_{1:t},x)}.$ (8)
This is essentially a variant of Zhong et al. (2025), where the token values within an action sequence are summed and averaged at each timestep. Please refer to Zhong et al. (2025) for proof details.
@@ 4 Experiments
We evaluate iStar across a variety of agentic tasks to demonstrate: (1) its effectiveness in training LLM agents for long-horizon reasoning and acting; (2) high sample efficiency and training stability given implicit step rewards; (3) improved exploration efficiency evidenced by increased rewards and fewer steps; and (4) the core components of iStar for credit assignment in agentic RL. Training details can be found in Appendix C.1.
@@
We evaluate LLM agents in three challenging environments: (1) WebShop (Yao et al., 2022), a text-based web environment where the agent interacts with a HTML-based website to search, nevigate, and purchase an item given an user instruction, requiring multi-step decision making; (2) VisualSokoban (Schrader, 2018) with $6\times 6$ size, a puzzle game where the agent has to push all boxes on targets, requiring spatial reasoning and long-term planning over both visual and textual inputs; (3) SOTOPIA (Zhou et al., 2024), an open-ended social interaction environment where the agent interacts with another LLM agent given a social scenario, role profiles and private goals, requiring reasoning over the other agent’s real-time strategies. During training, we use scenarios from SOTOPIA-$\pi$ (Wang et al., 2024b). See Appendix B for more details of the environments.
@@
We compare our method against a range of competitive baselines: (1) prompting LLMs specialized in general-purpose reasoning: GPT-5 (OpenAI, 2025b), Gemini-2.5-Pro (Google, 2024), DeepSeek-R1 (Guo et al., 2025), and Claude-Sonnet-4-Thinking (Anthropic, 2025); (2) vanilla RL methods that only use outcome rewards: PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), RLOO (Ahmadian et al., 2024) and REINFORCE++ (Hu et al., 2025); (3) a recent single-turn RL method: PRIME (Cui et al., 2025) that introduces token-level process rewards for policy learning; and (4) a recent agentic RL algorithm: GiGPO (Feng et al., 2025) that computes step-level advantages via same-state grouping.
@@
For WebShop and VisualSokoban, we adopt Success Rate and Score (only for WebShop) as the evaluation metrics following Feng et al. (2025). These metrics are computed over validation instances and select the best score for comparison. For SOTOPIA, we report goal completion score ranging from 0 to 10, which is evaluated by GPT-4o as a proxy for human judgement following Zhou et al. (2024). We set the tempature to 0 for the LLM judge. Refer to Appendix C.2 for evaluation prompts used in each environment.
Table 1: Performance on WebShop and VisualSokoban. Qwen2.5-7B-Instruct and Qwen2.5-VL-7B-Instruct serve the base models for WebShop and VisualSokoban, respectively. Note that Deepseek-R1 and PPO training do not currently support multi-modal scenarios, and PRIME is only applicable to tasks with binary outcome rewards. Results are averaged over three random seeds.
@@
+ RLOO w/ iStar $\textbf{86.5}\pm 2.8$ $\textbf{93.6}\pm 1.0$ $\textbf{91.7}\pm 1.2$
4.1 Main Results
Performance on benchmarks.
Table 1 showcases iStar’s superior performance over baselines in WebShop and VisualSokoban, with notable gains on the latter, where RL algorithms struggle with irreversible mistakes and limited foresight. Similar trends occur with much smaller base models as demonstrated in Table 4. Specifically, our method achieves state-of-the-art performance and surpasses recent multi-turn RL baseline GiGPO (Feng et al., 2025) by enabling finer-grained credit assignment, distinguishing good and bad actions with implicit rewards for each step than relying on same-state grouping. Our method also outperforms recent single-turn RL baseline PRIME (Cui et al., 2025) that uses token-level process rewards, which provides overly fine-grained rewards that complicate policy training in multi-turn RL with high-variance (see Figure 4(a)-(b)). In SOTOPIA, where GiGPO and PRIME are inapplicable due to open-ended state space and unverifiable rewards, Table 2 shows that iStar still achieves state-of-the-art performance. Particularly, compared to vanilla RL baselines, our method improves goal completion in hard social scenarios by 14% ($7.92\rightarrow 8.06$) in self-chat, and up to 48% ($6.68\rightarrow 7.16$) increase when chatting with GPT-4o. This demonstrates the generalizability of our method to a wide variety of interactive environments.
@@ 4.1 Main Results
+ GRPO w/ iStar $\textbf{7.11}\pm 0.19$ $\textbf{8.42}\pm 0.03$ $\textbf{6.76}\pm 0.18$ $\textbf{8.36}\pm 0.03$
@@
+ GRPO w/ iStar $\textbf{8.06}\pm 0.11$ $\textbf{9.20}\pm 0.03$ $\textbf{7.16}\pm 0.14$ $\textbf{8.45}\pm 0.03$
iStar with different vanilla RL algorithms.
Since our method is compatible with various RL methods, we also evaluate it using different RL algorithms, including RLOO (Ahmadian et al., 2024), REINFORCE++ (Hu et al., 2025) and GRPO (Shao et al., 2024), and compare each to its vanilla version that only uses outcome rewards. As shown in Figure 3, iStar consistently improves vanilla RL methods by integrating implicit step rewards into multi-turn RL to improve credit assignment in long-horizons. For example, iStar with RLOO obtains substantial gains of 6.3% in success rate on both WebShop and VisualSokoban. Similar trends occur with REINFORCE++ and GRPO, demonstrating the robustness of our method to diverse RL algorithms and environments.
Figure 3: Performance comparison of iStar with different vanilla RL algorithms. Qwen2.5-7B-Instruct and Qwen2.5-VL-7B-Instruct serve the base models for WebShop and VisualSokoban, respectively. Results are reported using one seed.
@@ 4.2 In-depth Analysis
Figure 4 illustrates that compared to baselines, iStar achieves faster improvement and higher final performance in validation metrics during RL training, demonstrating superior sample efficiency. In Figure 4(a), our method achieves the score of vanilla RLOO in WebShop in just 105 steps, around $2\times$ improvement in training efficiency. By 165 steps, we reach the highest score of 94.7%, showcasing significant improvements in both efficiency and performance. Notably, while PRIME exhibits comparable early-stage performance in WebShop, its growth stagnates and experiences sharp fluctuations. This is because overly fine-grained process rewards in token-level will complicate policy learning in multi-turn interactions, which usually involve much longer sequences than single-turn tasks. In contrast, iStar continues to improve consistently, suggesting that our implicit step rewards provide dense feedback for exploration while being sufficiently coarse to reduce variance for stable training.
Figure 4: Validation performance during multi-turn RL in WebShop and VisualSokoban. Note that PRIME can only be applied to tasks with binary outcome rewards.
Figure 5: Training dynamics of iStar in WebShop and VisualSokoban. Left: Dynamics of the episode and implicit step rewards during RL training by our method. Right: The episode length versus training step compared to baselines.
Exploration efficiency.
To demonstrate the implicit step rewards provide useful guidance for policy learning, we visualize the dynamics of both step- and episode-level rewards in Figure 5(a) and (b). In particular, the implicit step reward improves very early (especially in VisualSokoban) and then the episode reward follows, indicating that our method first captures good local action heuristics and then composes them into higher-return trajectories. As a result, it also reduces unnecessary actions during multi-turn interactions, leading to shorter episode lengths. As shown in Figure 5(c) and (d), episode lengths decrease without compromising task success, as evidenced by the consistent increase in episode rewards.
Table 3: Ablation studies on core components of iStar. “RLOO”: only outcome rewards are used to compute advantages for policy updates. “w/ ground-truth process rewards”: use raw step rewards provided by VisualSokoban to calculate step-level advantages. “w/ merged rewards”: implicit step rewards are added directly to outcome rewards before advantage computation. “w/ token-level process rewards”: implicit step rewards are computed in token level rather than step level. Results are reported using one seed.
@@
w/ ground-truth process rewards - - 87.5
@@
w/ token-level process rewards 82.0 90.0 89.1
w/ iStar 89.1 94.7 93.0
@@ 4.3 Ablation Studies
Table 3 presents results of ablation experiments to validate the necessity of key components of our method for effective credit assignment. First, raw step penalties provided in VisualSokoban show limited improvement over vanilla RL, suggesting that the implicit step rewards learned by iStar are superior credit signals for policy learning. Second, merging the implicit step rewards into episode rewards obtains gains over vanilla RL but the improvement are modest compared to our method. This indicates that we should not only reward intermediate actions but also gate credit by final task success to prevent speculative reward exploitation. Therefore, combining signals at the advantage level is crucial to credit assignment in long-horizons. Third, learning token-level process rewards is sub-optimal to multi-turn RL, suggesting that overly fine-grained rewards may introduce noise and thus increase difficulty of policy learning.
@@ 5 Related Work
PRMs has been widely explored in single-turn tasks, such as mathematical reasoning and single-shot code generation. In these settings, PRMs are used to score intermediate steps for test-time search or reranking (Lightman et al., 2024; Wang et al., 2024a; Mahan et al., 2024), or for online-RL (Dou et al., 2024; Setlur et al., 2025; Zha et al., 2025). Moving to dynamic, interactive tasks, process rewards are usually constructed in three ways: (1) handcrafted step rewards assigned to tool execution (Zeng et al., 2025) or meta-reasoning tags (Zhang et al., 2025b); (2) GRMs (Liu et al., 2025; Zou et al., 2025) that label step quality; (3) and implicit PRMs (Yuan et al., 2025) that produce token-level process rewards. However, manually-designed or judge-based step labels are costly and biased, suffering from reward hacking. Learning step Q-values (Choudhury, 2025) can reduce the bias but a fixed PRM may poorly estimate Q-values for unseen actions during inference. PRIME(Cui et al., 2025) partially addresses these issues by jointly training an implicit PRM with the generator. However, with overly fine-grained rewards in token-level, it introduces noises and distabilizes training in multi-turn RL. Another issue with PRIME is that it applies a cross-entropy loss to optimize the implicit PRM, which is only applicable to tasks with binary outcome rewards. Instead of learning a PRM, Feng et al. (2025) addresses credit assignment by computing step-level advantages via same-state grouping. While effective in tasks with finite state-action space, it relies on exact state overlaps and cannot be generalized to open-ended language environments where same state is rare. In contrast, our method learns implicit rewards at the step-level through a trajectory-level DPO objective, offering label-efficient step rewards with low-variance and showing robustness and generalization to (un)verifiable rewards even in open-ended environments.
6 Conclusion
We propose iStar, a general credit-assignment strategy for LLM agents. In particular, we alternatively optimize an implicit PRM with the policy model, which provides implicit step rewards to guide policy learning and results in a self-reinforcing training loop. to optimize the implicit PRM, we propose a trajectory-level DPO objective that is guaranteed to be equivalent to a BT model with a step-wise reward function. In practice, our method can be plugged into a variety of RL algorithms and generalized to (un)verifiable rewards even in open-ended environments. Empirical results show that iStar achieves averaged 86.5% success and 93.6% score on WebShop, and reaches 91.7% success on VisualSokoban. In SOTOPIA, our method improves goal completion in hard scenarios by up to 14% in self-chat and 48% against GPT-4o. In the future, our method can also be validated in math problems or code generation to provide implicit step rewards for intermediate CoT steps. It can also be applied to test-time scaling for search guidance. Regarding the limitations, we currently separate the implcit PRM from the policy model during training, which however, can be a unified model trained with different objectives to reduce computation memory and potentially improve representation sharing. Additionally, in SOTOPIA, our implicit PRM is trained to only predict goal-completion preferences, while future work could be extended to multi-objective implicit PRMs.